### v1

# ScaffoldGPT: A Scaffold-based Large Language Model for Drug Improvement

Abstract
Drug optimization has become increasingly crucial in light of fast-mutating virus strains and drug-resistant cancer cells. Nevertheless, it remains challenging as it necessitates retaining the beneficial properties of the original drug while simultaneously enhancing desired attributes beyond its scope. In this work, we aim to tackle this challenge by introducing ScaffoldGPT, a novel Large Language Model (LLM) designed for drug optimization based on molecular scaffolds. Our work comprises three key components: (1) A three-stage drug optimization approach that integrates pretraining, finetuning, and decoding optimization. (2) A uniquely designed two-phase incremental training approach for pre-training the drug optimization LLM-based generator on molecule scaffold with enhanced performance. (3) A token-level decoding optimization strategy, Top-N, that enabling controlled, reward-guided generation using pretrained/finetuned LLMs. Finally, by conducting a comprehensive evaluation on COVID and cancer benchmarks, we demonstrate that ScaffoldGPT outperforms the competing baselines in drug optimization benchmarks, while excelling in preserving the original functional scaffold and enhancing desired properties.
Machine Learning, ICML
[…]
1 Introduction
The rise of rapidly mutating virus strains [14], exemplified by those related to SARS-CoV-2 [49], along with drug-resistant cancer cells [27], has heightened the urgency and interest in accelerating the development of effective treatments. However, traditional De Novo drug discovery processes are prohibitively expensive, often costing from hundreds of millions to billions of dollars [5], due to their extensive and resource-demanding nature. Despite considerable efforts, the success rate of drug discovery remains low, with many candidates failing early in the development process. This has led to an increased focus on drug repurposing, which leverages existing FDA-approved drugs for new therapeutic uses rather than creating new drugs from scratch. While drug repurposing has seen some success [34], its effectiveness is often constrained because drugs are typically developed with a high specificity for a particular disease.
Drug improvement seeks to overcome the limitations inherent in both De Novo drug discovery and drug repurposing by enhancing an existing FDA-approved drug. Drug improvement is becoming an increasingly vital field, yet it remains relatively underexplored compared to drug discovery and repurposing efforts. DrugImprover [22] has started to clearly define the drug optimization problem by using Tanimoto similarity. Additionally, the DrugImprover framework contributes to the drug improvement domain in three key aspects: 1) a detailed workflow for drug optimization, 2) it introduced an Advantage-alignment Policy Optimization (APO) reinforcement learning (RL) algorithm to enhance the multi-objective generative model for drug improvement, and 3) includes an extensive dataset featuring 1 million ligands and their OEDOCK scores for five proteins related to human cancer cells and 24 high-affinity binding sites on the SARS-CoV-2 protein 3CLPro (PDB ID: 7BQY). The DrugImprover framework features a pretrained generative model that is subsequently refined using the APO reinforcement learning algorithm to ensure the molecules produced align with new objectives. Although DrugImprover has demonstrated encouraging outcomes, its effectiveness is limited due to its dependence on the less complex LSTM network architecture, which might lead to limited scalability and capacity, contextual understanding.
On the other hand, LLMs have shown exceptional performance in diverse areas, from natural language understanding and generation exemplified by ChatGPT [48, 32], to text-to-video conversion as seen in SORA [25], and in programming and code generation through platforms like PG-TD [50] and Copilot [30]. Nevertheless, in the field of drug discovery, tools such as DrugGPT [21] and ChemGPT [9], despite making some initial strides, have not yet achieved performance on par with their counterparts in other domains. The field of drug discovery remains in expectation of breakthroughs similar to the ones achieved by LLMs in other domains.
Several challenges have hindered the impact of large language models (LLMs) on drug design: Firstly, the molecules created need to satisfy multiple criteria such as solubility and synthesizability, and must also secure a high docking score against a specific target site. However, existing drug discovery LLMs generally only undergo pretraining with molecules and do not focus on enhancing multiple attributes simultaneously. Secondly, given that drugs sharing similar chemical structures tend to exhibit comparable biological or chemical effects [2], it is crucial to optimize the drug while preserving the beneficial chemical structure of the original molecule. Lastly, the current methodologies in drug discovery LLMs primarily focus on maximizing likelihood during the decoding phase, rather than customizing the optimization process to meet specific goals.
REINVENT 4 [15, 16, 26] achieves the state of art performance in LLM-based drug optimization by using the Transformer model and conducting experiments with randomly selected ligand pairs that maintain constrained Tanimoto Similarity. While this method successfully produces ligands with high Tanimoto Similarity, it encounters difficulties in consistently improving the properties of the original drug and tends to simply maximize the likelihood of molecule generation.
Our contributions. In this work, we propose ScaffoldGPT, a novel scaffold-based large language model with three-stage optimization process for drug improvement. This optimization process is designed to enhance existing drugs to rapidly evolving virus variants and cancer cells, overcoming the limitation of earlier drug improvement efforts. The contributions of ScaffoldGPT are summarized as follows:
$\bullet$ A framework for drug improvement using LLMs that involves a three-step optimization process. The underlying motivation is that each stage is complementary to each other, enhances the performance of the preceding one. Furthermore, we have conducted ablation studies to illustrate the performance gains achieved at each step.
$\bullet$ A scaffolds-based LLM with a novel two-phase incremental training specifically designed for drug optimization. The motivation for incremental training is to conduct local optimization before embarking on global optimization by breaking down the whole training corpus into knowledge pieces in an incremental order. We also conducted an ablation study in table 3 to demonstrate the effectiveness of the incremental training.
$\bullet$ A novel token-level decoding optimization strategy, Top-N, utilizes pretrained LLMs to enable controlled, reward-guided generation that aligns with targeted objectives.
$\bullet$ Through extensive experiments and ablation studies on real-world viral and cancer-related benchmarks, we demonstrate that ScaffoldGPT outperforms the competing baselines and reliably improves upon existing molecules/drugs in terms of desired targeted objectives while preserving original scaffold, resulting in superior drug candidates.
2 Related Work
2.1 Reinforcement Learning and RL finetuning
Reinforcement learning [41] has become a fundamental strategy in drug design [3, 13, 29, 31, 33, 39, 42, 45, 51, 52], focusing on optimizing rewards that aggregate predicted property scores from various pharmaceutical predictors. Traditional reinforcement learning methods in drug discovery often neglected molecular structure constraints, leading to significant structural changes and unsynthesizable compounds. Our research differs by refining existing drugs to enhance their attributes without redesigning them from scratch and using reinforcement learning to improve a pre-trained large language model generator, rather than starting anew. We also incorporate methods like Advantage-aligned Policy Optimization (APO) [22], which assigns rewards based on advantage preference over the original molecule, to fine-tune a Transformer model, ensuring it aligns with multiple pharmaceutical objectives while preserving molecular structure. This approach, which includes controllable decoding, refines the model beyond traditional reinforcement learning fine-tuning stages.
2.2 Planning with LLMs
Several studies have leveraged planning algorithms to enhance text outputs for a variety of NLP tasks. These include approaches like beam search optimization, machine translation improvements [37, 20, 4]. The PG-TD [50] method is tailored for code generation using a singular reward function, whereas ERP [23] introduces a novel concept by considering the certainty of each generated token along with an e-step forward entropy measurement to gauge potential outcomes. It has been shown that ERP effectively balances exploration and exploitation within molecular structures, leading to the discovery of high-reward molecules. Unlike previous studies that focus solely on planning with pre-trained large language models, our approach incorporates a novel decoding optimization as a critical final step in the algorithm. Moreover, our focus is on optimizing an existing drug rather than creating a new one from the ground up.
2.3 Drug Optimization
Recent efforts in drug optimization have primarily focused on a limited array of drug properties while often disregarding the docking score, an essential metric for evaluating structural compatibility with a target [52, 6]. DrugEx v3 [24] seeks to resolve this by using 3D molecular graphs that encapsulate more comprehensive data such as chemical valence rules. Nevertheless, this method’s complexity poses challenges in generating molecules that closely resemble the originals, thus leading to lower similarity. This reduced similarity could account for the diminished efficacy of graph-based methods, as deviations from the original molecular structures result in the loss of vital chemical properties. Conversely, our approach manages to preserve a significant level of similarity. DrugImprover [22] effectively begins to redefine the drug optimization challenge by employing reinforcement learning with a mix of multiple objectives as rewards. Still, it employs an LSTM in its generative model, which faces limitations in scalability, capacity, and understanding context. In contrast, our method utilizes a Transformer as the primary generative model, enhanced further with Advantage-aligned Policy Optimization (APO) for exploratory purposes and an optimized decoder for superior performance.
2.4 Large Language Model for Drug Discovery
Large language models like MolGPT [1] and ChemGPT [9] have been utilized in molecule generation and drug discovery [1, 36, 46], using formats like SMILES [47] and SELFIES [18] to standardize molecular representation. These models have shown promise in drug design, outperforming traditional methods in some predictive capacities. However, as noted by  Murakumo et al. [28], using pre-trained large language models typically results in minor molecular modifications and serves primarily supportive roles in the design process. Our research adopts the SMILES format and combines the strengths of large language models with reinforcement learning. Unlike previous efforts that relied solely on LLMs, our approach uses RL fine-tuning to align with multiple pharmaceutical objectives and employs controlled decoding to guide LLMs in generating more effective molecular structures. Further developments include transformer-based models like REINVENT 4  [15, 16, 26], which mostly focuses on pretraining. While pretraining helps in generating molecules similar to those in the training dataset, it also inherently restricts the exploration scope due to biases in the training data. REINVENT4 uses the original molecule as input, generates molecules that are very similar to the original ones. This likely contributes to its relatively poor performance in other metrics, as the generated molecules exhibit minimal changes in chemical properties. In contrast, our proposed method, which uses SMILES and scaffolds as prompts, achieves a well-balanced trade-off between diversity and similarity in the generated and original molecules, potentially leading to improved performance.
3 Preliminaries
In the following sections, we detail MDP, LLM and drug discovery, complete with their mathematical notations, and integrate them within the framework of Markov decision processes.
Markov decision processes. Let us define a finite-horizon Markov decision process (MDP) [35] $\mathcal{M}_{0}=\langle\mathcal{S},\mathcal{A},T,\mathcal{P},R\rangle$. In this context, $\mathcal{S}$ represents a finite set of states, while $\mathcal{A}$ comprises a finite set of actions. The term $T$ denotes the planning horizon. The function $\mathcal{P}$, defined as $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}^{\prime}$, describes the deterministic transition dynamics that combine a state $s$ with an action $a$, with an episode concluding once the agent executes the termination action. Additionally, the reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scores exclusively to complete molecules, assigning a reward of 0 to partial molecules. The effectiveness of a policy can be evaluated using the $Q$-value function, denoted as $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and defined by the following equation: $Q^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{t=0}^{T}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right],$ where the expectation is based on the trajectory determined by the policy $\pi$. The associated value function is given by: $V^{\pi}(s):=\mathbb{E}_{a\sim\pi(\cdot|s)}\left[Q^{\pi}(s,a)\right]$.
[…]
$\mathcal{Y}_{T}:=\left\{\text{[BOS]}\circ\mathbf{v}\circ\text{[EOS]}~{}|~{}\mathbf{v}\in\mathcal{V}^{*}\right\},$ (1)
[…]
Drug Optimization. Given an initial drug candidate $X=\left({x_{1},\cdots,x_{T}}\right)$ and a drug optimization policy $\pi_{\theta}$, the goal in drug optimization is to find the optimal policy $\pi_{\theta^{*}}$ that maximize the following objective:
$\pi_{\theta^{*}}=\mathop{\rm arg\,max}_{\pi_{\theta}}\mathbb{E}_{X\sim{d_{0}}}\!\left[R(Y)-R\left({X}\right)|\theta,X\right],$ (2)
where $Y=\pi_{\theta}\left({\cdot{\,|\,}X}\right),Y_{1:T}=\left({y_{1},\ldots,y_{t},\ldots,y_{T}}\right),y_{t}\in\mathcal{V}.$
Limitation of previous works: DrugImprover, which utilizes LSTM networks, has limitations in scalability, capacity, and contextual understanding, especially when compared to versions that use Large Language Models. The current state of the art, REINVENT 4, employs the Transformer architecture; however, it mainly focuses on pretraining with constrained similarity, which restricts its capability to explore molecular spaces that might offer high rewards beyond its training set. In this work, we address these limitations by proposing ScaffoldGPT.
4 The ScaffoldGPT Algorithm
Algorithm 1 ScaffoldGPT
[…]
Figure 1: Two-phase incremental training of ScaffoldGPT. The first phase concentrates on recognizing the molecule, while the second phase builds connections between its scaffold and the original molecule.
Stage 1. Pretrain a LLM-based generator. Let us note the LLM-based generator policy as $\pi_{\theta}$, which computes the probability $p$ of the occurrence of the $t^{th}$ token in a target molecule $Y$. It takes into account all preceding tokens $\mathbf{y}_{<t}=\left[y_{1},...,y_{t-1}\right]$ in the target, as well as the scaffold compound $S$, which is noted as $\pi_{\theta}\left({y_{t}{\,|\,}\mathbf{y}_{<t},S}\right)=p\left({y_{t}|\mathbf{y}_{<t},S}\right).$ The parameters $\theta$ of the generator policy $\pi_{\theta}$ are trained using the training corpus set through the minimization of the negative log-likelihood (NLL) for the complete SMILES strings across the entire set. This process is described as follows:
$NLL$ $=-\log{P}\left({Y|S}\right)=-\sum_{t=1}^{T}\log{P}\left({y_{t}|y_{t-1},...,y_{1},S}\right)$
$=-\sum_{t=1}^{T}\log\pi_{\theta}\left({y_{t}|y_{1:t-1},S}\right),$ (3)
where $T$ signifies the total number of tokens related to $Y$. The NLL quantifies the probability of converting a specific scaffold into a designated target molecule.
In this project, we employ pre-training to harness large quantities of unlabeled text to construct a basic foundation model of language understanding. This foundation model can subsequently be fine-tuned and tailored to meet various specialized goals. In this work, we propose a novel framework for pre-training a Transformer and linking scaffolds with complete molecules, based on a SMILES (Simplified Molecular Input Line Entry System) [47] string representation of the molecule. Here, we define $\mathcal{S}$ as the scaffold space and $\mathcal{Y}$ as the target molecule space. With $\mathcal{P}=\left\{\left({s,y}\right)|s,y\in\mathcal{S}\times\mathcal{Y}\right\}$, we denote the set of scaffold and molecular pairs from $\mathcal{S}$ and $\mathcal{Y}$. In this notation, $s$ represents the scaffold of the target molecule, and $y$ is the corresponding target molecule. We initially pre-trained our tokenizer using the Byte Pair Encoding (BPE) method [11, 38]. Building on the pre-trained BPE tokenizer, we propose a two-phase incremental training approach, as illustrated in Fig. 1, to notably enhance the model’s ability to improve the validity of inferring the target molecule from its scaffold.
Incremental training.
The rationale for incremental training is to conduct local optimization before embarking on global optimization. Therefore, we divide our training into two phases. In the first phase, we focus on training a large language model (LLM) exclusively for molecules using Causal Language Modeling (CLM). CLM utilizes an autoregressive method where the model is trained to predict the next token in a sequence by considering only the tokens that precede it. The phase 1 corpus is designed as follows:
$C_{\text{Phase 1}}=\left\{[BOS],\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},$ (4)
In the second phase, building upon the success of the LLM model developed in phase 1, which demonstrated high accuracy in molecular generation, we advance the training by focusing on pairs of scaffolds and molecules using CLM. The phase 2 corpus is designed as follows:
$C_{\text{Phase 2}}=\left\{\left\langle S\right\rangle,\underbrace{s_{1},\cdots,s_{T}}_{\text{source scaffold S}},\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}}\right\},$ (5)
Consequently, the model can generate the appropriate molecule when given a scaffold. However, since a single scaffold may correspond to multiple molecules, we further refine the LLM-based generator policy, $\pi_{\theta}$, to target specific outcomes by applying reinforcement learning finetuning in the next stage.
Stage 2. RL finetuning. Fine-tuning a generative model is crucial for producing outcomes that meet specific objectives. In this study, we use the Advantage-alignment Policy Optimization (APO) [22] algorithm to fine-tune the pretrained LLM-based generator policy $\pi_{\theta}$. This approach steers the model from a given scaffold towards the targeted molecule, simultaneously improving multiple properties.
In this work, we adopt the define of reward function from Liu et al. [23], and regarded each pharmaceutical property as a critic $C$ and got an ensemble critics $\mathbf{C}$ as follows:
$[{C}^{\text{Druglikeness}},{C}^{\text{Solubility}},{C}^{\text{Synthesizability}},{{C}^{\text{Docking}}},{C}^{\text{{Tanimoto}}}],$
where each critic ${C}:Y\rightarrow\mathbb{R}$ acts as a distinct evaluator for a specific pharmaceutical attribute. We built the reward function as follow:
${R^{\text{norm}}\left({Y}\right)}:={R^{\text{norm}}\left({Y|S}\right)}$ (6)
$=\lambda\cdot\text{Norm}\left({{C}^{\text{Tanimoto}}\left({S,Y}\right)}\right)+\sum_{i=0}^{|\mathbf{C}|-1}\lambda\cdot\text{Norm}\left({{C}_{i}{\left({Y}\right)}}\right),$
where Norm is employed to standardize diverse attributes to a consistent scale. In this instance, Norm refers to the process of min-max normalization, which is used to adjust the attributes so they fit within the $\left[0,1\right]$ range.
In this work, we use BON [12] (Best of N) search to estimate the reward for a prompt (partial molecule). BON can be formulated as the following:
BON $\left({\mathbf{y}_{<i},N,R}\right)|_{S,p,k}=\max_{\mathbf{Y}_{j}\in\left\{\mathbf{Y}_{1},\cdots,\mathbf{Y}_{N}\right\}}R\left({\mathbf{Y}_{j}}\right),$ (7)
where $\text{where }Y_{j}=\left[\mathbf{y}_{<i},y_{i},\cdots,y_{T}\right]_{j},$
$\text{and }y_{i}\sim\textsc{Top-PK}\left({\mathbf{y}_{<i},p,k}\right)|_{S}.$
where Top-PK [23] defined as follows:
[…]
 $\text{~{}~{}~{}~{}~{}~{}~{}}j=\min\left\{\mathop{\rm arg\,min}_{j^{\prime}}\sum_{i=1}^{j^{\prime}}\pi_{\theta}\left({y_{i}|S,\mathbf{y}_{<i}}\right)\geq p,k\right\},$
 $\text{~{}~{}~{}~{}~{}~{}~{}}\text{and }\pi_{\theta}\left({y_{g}|S,\mathbf{y}_{<i}}\right)>\pi_{\theta}\left({y_{h}|S,\mathbf{y}_{<i}}\right),\text{if }g<h,$
where $p\in(0,1]$ represents the maximum cumulative probability, and $k$ denotes the maximum number of candidates for the next tokens.
Target Algorithm Validity $\uparrow$ Avg Norm Reward $\uparrow$ Avg Top 10 % Norm Reward $\uparrow$ Docking  $\downarrow$ Druglikeliness  $\uparrow$ Synthesizability  $\downarrow$ Solubility  $\uparrow$ Similarity $\uparrow$
3CLPro Original - 0.533 0.689 -8.698 0.682 3.920 2.471 -
(PDBID: MMP [26] 0.995 $\pm$ 0.001 0.628 $\pm$ 0.001 0.718 $\pm$ $0$8.259 $\pm$ 0.004 0.691 $\pm$ 0.001 2.682 $\pm$ 0.004 3.109 $\pm$ 0.020 0.862 $\pm$ 0.000
7BQY) Similarity ($\geq$ 0.5) [26] 0.995 $\pm$ 0.001 0.615 $\pm$ 0.000 0.706 $\pm$ 0.001 -8.165 $\pm$ 0.024 0.697 $\pm$ 0.004 2.621 $\pm$ 0.006 3.180 $\pm$ 0.029 0.782 $\pm$ 0.001
Similarity ([0.5, 0.7)]) [26] 0.995 $\pm$ 0.001 0.612 $\pm$ 0.001 0.701 $\pm$ 0.001 -8.187 $\pm$ 0.010 0.691 $\pm$ 0.001 2.611 $\pm$ 0.009 3.240 $\pm$ 0.014 0.756 $\pm$ 0.003
Similarity ($\geq$ 0.7) [26] 0.995 $\pm$ 0.001 0.628 $\pm$ 0.001 0.718 $\pm$ 0.001 -8.214 $\pm$ 0.002 0.691 $\pm$ 0.002 2.717 $\pm$ 0.002 3.080 $\pm$ 0.016 0.881 $\pm$ 0.002
Scaffold [26] 0.995 $\pm$ 0.001 0.602 $\pm$ 0.001 0.703 $\pm$ 0.002 -8.116 $\pm$ 0.002 0.695 $\pm$ 0.001 2.728 $\pm$ 0.008 2.968 $\pm$ 0.038 0.776 $\pm$ 0.001
Scaffold Generic [26] 0.994 $\pm$ 0.001 0.617 $\pm$ 0.001 0.710 $\pm$ 0.002 -8.179 $\pm$ 0.012 0.701 $\pm$ 0.000 2.645 $\pm$ 0.008 3.090 $\pm$ 0.029 0.801 $\pm$ 0.000
DrugImprover [22] 0.884 $\pm$ 0.005 0.432 $\pm$ 0.002 0.493 $\pm$ 0.005 -6.726 $\pm$ 0.007 0.506 $\pm$ 0.002 1.306 $\pm$ 0.010 2.057 $\pm$ 0.011 0.531 $\pm$ 0.002
Molsearch [40] 1.000 $\pm$ 0.001 0.616 $\pm$ 0.001 0.726 $\pm$ 0.002 -8.855 $\pm$ 0.040 0.686 $\pm$ 0.001 3.105 $\pm$ 0.006 2.452 $\pm$ 0.008 0.969 $\pm$ 0.001
MIMOSA [10] 0.985 $\pm$ 0.008 0.622 $\pm$ 0.001 0.734 $\pm$ 0.002 -8.800 $\pm$ 0.015 0.677 $\pm$ 0.004 3.105 $\pm$ 0.008 2.711 $\pm$ 0.010 0.959 $\pm$ 0.001
DrugEx v3 [24] 1.000 $\pm$ 0.001 0.524 $\pm$ 0.001 0.613 $\pm$ 0.001 -8.089 $\pm$ 0.013 0.583 $\pm$ 0.002 3.095 $\pm$ 0.005 3.932 $\pm$ 0.008 0.495 $\pm$ 0.001
ScaffoldGPT (w/o APO & Top-N) 0.951 $\pm$ 0.004 0.587 $\pm$ 0.004 0.693 $\pm$ 0.004 -8.238 $\pm$ 0.101 0.659 $\pm$ 0.014 2.865 $\pm$ 0.038 2.999 $\pm$ 0.163 0.754 $\pm$ 0.005
ScaffoldGPT (w/o Top-N) 0.857 $\pm$ 0.061 0.627 $\pm$ 0.009 0.717 $\pm$ 0.004 -8.583 $\pm$ 0.075 0.727 $\pm$ 0.019 2.566 $\pm$ 0.088 3.388 $\pm$ 0.095 0.717 $\pm$ 0.028
ScaffoldGPT (w/o APO) 0.998 $\pm$ 0.001 0.666 $\pm$ 0.000 0.740 $\pm$ 0.001 -9.312 $\pm$ 0.018 0.734 $\pm$ 0.002 2.698 $\pm$ 0.006 3.676 $\pm$ 0.006 0.813 $\pm$ 0.002
ScaffoldGPT 0.944 $\pm$ 0.094 0.675 $\pm$ 0.031 0.740 $\pm$ 0.015 -9.343 $\pm$ 0.440 0.746 $\pm$ 0.028 2.453 $\pm$ 0.154 3.913 $\pm$ 0.358 0.745 $\pm$ 0.032
RTCB Original - 0.536 0.698 -8.572 0.709 3.005 2.299 -
(PDBID: MMP [26] 0.998 $\pm$ 0.001 0.636 $\pm$ 0.000 0.731 $\pm$ 0.001 -8.465 $\pm$ 0.021 0.709 $\pm$ 0.001 2.599 $\pm$ 0.004 3.013 $\pm$ 0.013 0.845 $\pm$ 0.001
4DWQ) Similarity ($\geq$ 0.5) [26] 0.999 $\pm$ 0.001 0.626 $\pm$ 0.000 0.723 $\pm$ 0.001 -8.511 $\pm$ 0.012 0.713 $\pm$ 0.002 2.543 $\pm$ 0.002 3.082 $\pm$ 0.031 0.760 $\pm$ 0.000
Similarity ([0.5, 0.7)]) [26] 0.999 $\pm$ 0.001 0.622 $\pm$ 0.001 0.718 $\pm$ $0$8.486 $\pm$ 0.021 0.713 $\pm$ 0.003 2.542 $\pm$ 0.005 3.101 $\pm$ 0.005 0.740 $\pm$ 0.001
Similarity ($\geq$ 0.7) [26] 0.999 $\pm$ 0.001 0.639 $\pm$ 0.000 0.734 $\pm$ 0.001 -8.496 $\pm$ 0.009 0.718 $\pm$ 0.001 2.628 $\pm$ 0.001 2.868 $\pm$ 0.003 0.875 $\pm$ 0.002
Scaffold [26] 0.998 $\pm$ 0.001 0.609 $\pm$ 0.001 0.718 $\pm$ $0$8.508 $\pm$ 0.026 0.711 $\pm$ 0.000 2.627 $\pm$ 0.002 2.803 $\pm$ 0.010 0.735 $\pm$ 0.002
Scaffold Generic [26] 0.998 $\pm$ 0.001 0.625 $\pm$ 0.001 0.722 $\pm$ $0$8.544 $\pm$ 0.009 0.722 $\pm$ 0.002 2.551 $\pm$ 0.010 2.898 $\pm$ 0.005 0.768 $\pm$ 0.004
DrugImprover [22] 0.920 $\pm$ 0.008 0.478 $\pm$ 0.001 0.618 $\pm$ 0.002 -8.701 $\pm$ 0.037 0.486 $\pm$ 0.002 1.181 $\pm$ 0.010 2.026 $\pm$ 0.013 0.427 $\pm$ 0.001
Molsearch [40] 1.000 $\pm$ 0.001 0.625 $\pm$ 0.001 0.742 $\pm$ 0.001 -8.747 $\pm$ 0.009 0.719 $\pm$ 0.001 3.012 $\pm$ 0.004 2.273 $\pm$ 0.005 0.950 $\pm$ 0.001
MIMOSA [10] 0.989 $\pm$ 0.001 0.631 $\pm$ 0.001 0.749 $\pm$ 0.001 -8.972 $\pm$ 0.011 0.706 $\pm$ 0.003 3.080 $\pm$ 0.007 2.561 $\pm$ 0.008 0.945 $\pm$ 0.001
DrugEx v3 [24] 1.000 $\pm$ 0.001 0.592 $\pm$ 0.001 0.668 $\pm$ 0.001 -8.762 $\pm$ 0.010 0.583 $\pm$ 0.002 2.488 $\pm$ 0.005 5.827 $\pm$ 0.010 0.393 $\pm$ 0.001
ScaffoldGPT (w/o APO & Top-N) 0.956 $\pm$ 0.004 0.582 $\pm$ 0.007 0.700 $\pm$ 0.008 -8.214 $\pm$ 0.125 0.686 $\pm$ 0.017 2.788 $\pm$ 0.056 2.781 $\pm$ 0.214 0.707 $\pm$ 0.005
ScaffoldGPT (w/o Top-N) 0.811 $\pm$ 0.074 0.639 $\pm$ 0.004 0.723 $\pm$ 0.005 -8.808 $\pm$ 0.071 0.741 $\pm$ 0.013 2.521 $\pm$ 0.081 3.279 $\pm$ 0.067 0.730 $\pm$ 0.030
ScaffoldGPT (w/o APO) 0.997 $\pm$ 0.001 0.673 $\pm$ 0.001 0.755 $\pm$ 0.001 -9.659 $\pm$ 0.023 0.764 $\pm$ 0.001 2.606 $\pm$ 0.007 3.481 $\pm$ 0.027 0.773 $\pm$ 0.003
ScaffoldGPT 0.826 $\pm$ 0.100 0.682 $\pm$ 0.004 0.756 $\pm$ 0.003 -9.757 $\pm$ 0.057 0.765 $\pm$ 0.013 2.437 $\pm$ 0.059 3.582 $\pm$ 0.043 0.747 $\pm$ 0.026
Table 1: Main results. A comparison of eight baselines including Original, six baselines from REINVENT {MMP, Similarity ($\geq 0.5$), Similarity $\in[0.5,0.7)$, Similarity $\geq 0.7$, Scaffold, Scaffold Generic}, DrugImprover, Molsearch, MIMOSA and different versions of ScaffoldGPT on various objectives based on 3CLPro and RTCB datasets. The top two results are highlighted as 1st and 2nd. Results are reported for 5 experimental runs.
For each pair consisting of a scaffold and a molecule, we create 8 new molecules from the scaffold using Top-PK (8) and BON (7) method to select the best one (the one with the highest normalized reward) to serve as the foundation for the final reward calculation.
${R_{c}\left({Y|S}\right)}=R^{\text{norm}}{\left({Y}\right)}|_{S},~{}~{}{Y}\in\text{{{BON}}}\left({y_{0},N,R}\right)|_{S,p,k}.$ (9)
APO makes policy gradient based on the advantage preference [22], which is defined as
$R^{\text{AP}}\left({Y_{1:T},S}\right)={R_{c}\left({Y_{1:T}}\right)-R_{c}\left({S}\right)},$ (10)
and perform APO policy gradient with follows:
${g}=$ $\textstyle\mathbb{E}_{S\sim{\rho_{0}},{Y_{1:T}\sim\pi_{{\theta}}\left({\cdot|S}\right)}}\left[\nabla_{\theta}\log\pi_{{\theta}}\left({Y_{1:T}|S}\right)\cdot R^{\mathtt{AP}}\left({S,Y_{1:T}}\right)\right],$ (11)
Stage 3. Token-level Controllable decoding generation.
Ultimately, the current LLM-based decoder focuses mainly on maximizing likelihood, neglecting specific metrics of interest. This approach limits its effectiveness in optimizing objectives that diverge from those in its training set, particularly in generating desired molecules. In this study, we introduce controllable decoding after fine-tuning with APO. We present a new approach, Top-N, to direct the generation process towards enhancements in the optimization objective, as detailed below:
$Y^{\star}$$\sim\left\{[\mathbf{y}_{<i},y_{i},\cdots,y_{T}]\right\},$ (12)
[…]
 $\text{~{}}\mathcal{A}_{\mathbf{y}_{<i}}=\left\{y_{1},\ldots,y_{n}\right\},y_{i}\in\mathcal{V},|\mathcal{A}|\leq k,$
 and $\text{~{}and }R\left({\text{{{BON}}}\left({\mathbf{y}_{<i}\circ y_{g},N,R}\right)|_{S,p,k}}\right)$
$\geq R\left({\text{{{BON}}}\left({\mathbf{y}_{<i}\circ y_{h},N,R}\right)|_{S,p,k}}\right),\forall g<h,$
where $n\leq k$ denotes as top $N$ candidate of next tokens with regard to BON function.
Remark 4.1.
Top-N differs from Top-P, Top-K, and Top-PK in that it is measured based on maximum reward, whereas the others are measured based on maximum likelihood. Top-N is also distinct from BON in that BON is optimized at the sequence level, while Top-N is optimized at the token level.
5 EXPERIMENTS
5.1 Experimental configuration
The language model. We utilize GPT-2-like Transformers for causal language modeling, employing the standard 11M Drug-like Zinc dataset for training. Entries with empty scaffold SMILES are excluded, and we adopt a 90/10 split for training and validation, respectively. The training process is structured into three phases: pretraining, fine-tuning, and decoding optimization, as outlined in Algorithm ScaffoldGPT (See appendix for more details).
Original                                 Scaffold improved 1 improved 2 improved 3
Molecule [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE String Cc1ccc(O)c(Nc2nnc (-c3ccccc3)c(=O)[nH]2)c1 Cc1ccc(-c2nnc(Nc3ccc (C(C)C)cc3)[nH]c2=O)cc1 COc1ccc(-c2nnc(Nc3cccc (C)c3)[nH]c2=O)cc1 Cc1ccc(-c2nnc(Nc3cccc (C(C)C)c3)[nH]c2=O)cc1
Scaffold O=c1[nH]c(Nc2ccccc2)nnc1-c1ccccc1 same same same
Similarity ()($\uparrow$)( ↑ ) - 0.826 0.911 0.866
Docking ()($\downarrow$)( ↓ ) -10.031 -11.478 $\checkmark$ -11.474 $\checkmark$ -11.087 $\checkmark$
Druglikeness ()($\uparrow$)( ↑ ) 0.646 0.762 $\checkmark$ 0.774 $\checkmark$ 0.762 $\checkmark$
Synthesizability ()($\downarrow$)( ↓ ) 2.390 2.298 $\checkmark$ 2.257 $\checkmark$ 2.356 $\checkmark$
Solubility ()($\uparrow$)( ↑ ) 2.590 4.007 $\checkmark$ 2.893 $\checkmark$ 4.007 $\checkmark$
Avg Norm Reward ()($\uparrow$)( ↑ ) 0.618 0.759 $\checkmark$ 0.753 $\checkmark$ 0.754 $\checkmark$
Table 2: One optimization example from cancer benchmark. Every generated molecules retains the scaffold, with all desired properties improved compared to the original. $\checkmark$ indicates improved property.
Baselines. In this work, we employ baseline models including DrugImprover [22], which utilizes an LSTM-based generator with APO fine-tuning, Molsearch [40], a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization, MIMOSA [10], a sampling-based method leveraging graph-based molecular optimization and DrugEx v3 [24], which leverages transformer-based reinforcement learning for scaffold-guided drug optimization. Additionally, we incorporate the model proposed by He et al. [15, 16], Loeffler et al. [26], which trains a transformer to adhere to the Matched Molecular Pair (MMP) guidelines [17, 43]. Specifically, given a set $\{\{X,Y,Z\}\}$, where $X$ represents the source molecule, $Y$ denotes the target molecule, and $Z$ signifies the property change between $X$ and $Y$, the model learns a mapping from $\{X,Z\}\in\mathcal{X}\times\mathcal{Z}\implies Y\in\mathcal{Y}$ during training. Here, $\mathcal{X}\times\mathcal{Z}$ denotes the input space, and $\mathcal{Y}$ denotes the target space. They defined six different types of property changes for $Z$, including MMP for user-specified alterations, various similarity thresholds, and scaffold-based modifications where molecules share the same scaffold or a generic scaffold. More specifically,
  • MMP: there exists user-specified desirable property changes between molecule $X$ and $Y$.
  • Similarity $\geq 0.5$: the tanimoto similarity between molecule $X$ and $Y$ is larger than 0.5.
  • Similarity $\in[0.5,0.7)$: the tanimoto similarity of pair $\left({X,Y}\right)$ is between 0.5 and 0.7.
  • Similarity $\geq 0.7$: the tanimoto similarity between molecule $X$ and $Y$ is larger than 0.7.
  • Scaffold: molecule $X$ and $Y$ share same scaffold.
  • Scaffold generic: molecule $X$ and $Y$ share same generic scaffold.
All baseline models are fine-tuned on the cancer and COVID datasets following their respective fine-tuning methods.
Dataset. We employ, from the most recent Cancer and COVID dataset of Liu et al. [22], 1 million compounds from the ZINC15 dataset docked to the 3CLPro (PDB ID: 7BQY) protein associated with SARS-CoV-2 and the RTCB (PDB ID: 4DWQ) human cancer protein.
Critics and evaluation metric. In this study, we evaluate the efficacy of ScaffoldGPT in generating molecules with desirable attributes within the context of pharmaceutical drug discovery. We leverage the RDKit [19] chemoinformatics package and employ various performance metrics as follows: Validity: It measures if the generated SMILES is valid in syntax. Druglikeness: The druglikeness metric measures the likelihood of a molecule being a suitable candidate for drug development. Solubility: This metric assesses the likelihood of a molecule’s ability to mix with water, commonly referred to as the water-octanol partition coefficient (LogP). Synthetizability: This parameter quantifies the ease (score of 1) or difficulty (score of 10) associated with synthesizing a given molecule [7]. Docking Score: The docking score assesses the drug’s potential to bind and inhibit the target site. To enable efficient computation, we employ a docking surrogate model (See Appendix A.4) to output this score. Similarity: We use Tanimoto similarity to evaluate the similarity between original SMILES and generated SMILES. Average Top 10% Norm Reward: It is the average of the normalized reward of the top 10% of molecules based on their average normalized reward. Average Norm Reward: It is the average of the normalized values of the docking score, druglikeness, synthesizability, solubility, and similarity across all valid molecules. This is the most important metric.
5.2 Main results.
Table 1 shows that ScaffoldGPT surpasses DrugImprover and six different versions of REINVENT4 in performance measures for both virus-related and cancer-related proteins. Moreover, ScaffoldGPT exceeds the performance of all baseline methods and also demonstrates a decent level of Tanimoto similarity to the original drug, indicating that it preserves the advantageous features of the original drugs while improving desired properties.
Several key factors contribute to this superior performance. Although DrugImprover established a strong foundation for the drug optimization field, including a workflow and a reinforcement learning algorithm to align the generative model with multiple pharmaceutical objectives, ScaffoldGPT outshines DrugImprover in all benchmarks. This is because ScaffoldGPT employs a GPT-2-like Transformer as the basis of its generative model, whereas DrugImprover relies solely on LSTM. Consequently, the GPT-2 Transformer grants ScaffoldGPT enhanced scalability, capacity, and contextual understanding compared to DrugImprover.
In contrast to the current state-of-the-art approach, REINVENT4, which pre-trains a Transformer with constraints on Tanimoto similarity, their method falls short in achieving drug improvement as it overlooks the optimization of multiple pharmaceutical properties. Therefore, Table 1 reveals that although REINVENT4 achieved high similarity, the generated molecules often failed to surpass the original ones. ScaffoldGPT, on the other hand, employs the APO reinforcement learning algorithm to fine-tune the pre-trained generative model and utilizes the Top-N decoding optimization strategy. These approaches ensure improvements aligned with multiple pharmaceutical objectives and enable ScaffoldGPT to successfully enhance the original drug across various pharmaceutical properties while maintaining a high Tanimoto similarity.
5.3 Ablation studies.
1024 SMILES Details
1024 $>100$ length, $>50$ scaffold
Validity
One phase 0.49
Two phases 0.67
Table 3: Ablation study of one-phase vs two-phases
Model Diversity $\uparrow$ Validity $\uparrow$ Avg Norm Reward $\uparrow$ Avg Top 10% Norm Reward $\uparrow$ Docking $\downarrow$ Druglikeliness $\uparrow$ Synthesizability $\downarrow$ Solubility $\uparrow$ Similarity $\uparrow$
COVID
ScaffoldGPT (w/o Top-N) Top-K 0.988 0.947 0.645 0.727 -8.726 0.760 2.418 3.499 0.697
ScaffoldGPT (w/o Top-N) Top-P 0.988 0.938 0.642 0.722 -8.653 0.756 2.420 3.505 0.696
ScaffoldGPT (w/o Top-N) Top-PK 0.989 0.941 0.643 0.724 -8.667 0.759 2.407 3.506 0.692
ScaffoldGPT (Top-N) 0.965 0.944 0.675 0.740 -9.343 0.746 2.453 3.913 0.745
CANCER
ScaffoldGPT (w/o Top-N) Top-K 0.912 0.709 0.648 0.728 -8.944 0.756 2.456 3.258 0.730
ScaffoldGPT (w/o Top-N) Top-P 0.931 0.704 0.645 0.729 -8.907 0.756 2.466 3.226 0.723
ScaffoldGPT (w/o Top-N) Top-PK 0.926 0.719 0.645 0.727 -8.888 0.757 2.466 3.219 0.725
ScaffoldGPT (Top-N) 0.916 0.826 0.682 0.756 -9.757 0.765 2.437 3.582 0.747
Table 4: Ablation study of Top-N, Top-P, Top-K and Top-PK sampling strategies. The top result is highlighted as 1st. Top-N outperforms others in most of the metrics.
In this section, we present ablation studies that underscore the necessity and effectiveness of each component of ScaffoldGPT. These components complement each other, substantially enhancing overall performance.
Effectiveness of (two-phase) incremental training .
In two-phase incremental pretraining, the intuition behind the first phase lies in training on critical keywords as knowledge pieces, reinforcing the memory of these key terms, particularly in longer sequences. We conducted an ablation study comparing two-phase incremental training with one-phase training in Table 3. The results showed that the two-phase approach significantly improves validity compared to one-phase training, demonstrating the effectiveness of the incremental training method.
Effectiveness of APO Finetuning.
ScaffoldGPT adopts APO finetuning as the second step, following the completion of pretraining the LLM-based generator. Table 1 demonstrates the effectiveness of APO through two comparisons: ScaffoldGPT (w/o APO, Top-N) vs. ScaffoldGPT (w/o Top-N), which shows that after applying APO finetuning, performance improved on most properties. Additionally, ScaffoldGPT vs. ScaffoldGPT (w/o APO) validates the importance of the APO component. By applying APO on top of pretraining and Top-N decoding, performance improved. Both cases demonstrate the effectiveness of APO finetuning.
Effectiveness of Top-N decoding strategy.
ScaffoldGPT adopts the Top-N decoding strategy as the final step followed by APO finetuning. Table 1 demonstrates the effectiveness of Top-N through two comparisons: ScaffoldGPT (w/o APO, Top-N) vs. ScaffoldGPT (w/o APO), showing that after applying the Top-N decoding strategy on top of pretrained LLM, performance improved across most properties. Moreover, ScaffoldGPT vs. ScaffoldGPT (w/o Top-N) illustrates that after applying APO on top of pretraining and RL, performance still improves on multiple attributes, surpassing all baselines. Furthermore, by comparing ScaffoldGPT (w/o APO) and ScaffoldGPT (w/o Top-N), we observe that applying Top-N decoding alone enhances performance more than applying APO alone.
Ablation study of Top-N vs Top-P, Top-K and Top-PK strategies.
When removing Top-N component, we employ multinomial generation, where multinomial sampling randomly selects the next token from the entire vocabulary based on the model’s probability distribution. In the ablation study detailed in Table 4, we examined Top-N over Top-P, Top-K, and Top-PK with K=20 and P=0.95. The results indicate that Top-N surpasses the other strategies in most metrics.
Drug optimization illustration.
Finally, we provide three examples illustrating the effectiveness of ScaffoldGPT in improving upon the original molecule on the cancer benchmark, as shown in Table 2 (Refer to Appendix A.5 for the COVID benchmark). The results in Table 2 demonstrate that the drugs generated by ScaffoldGPT outperform the original drugs across all desired properties. Additionally, the comparison figure in Table 2 illustrates that the improved molecules preserve the original drug to a significant extent, with only minor changes highlighted in red. The results indicate that ScaffoldGPT effectively optimizes desired properties while preserving the beneficial properties of the original drug.
6 Conclusion
In this study, we introduce ScaffoldGPT, a novel framework named designed for drug optimization. It incorporates a unique scaffold-based LLM design, a three-step optimization process, a two-phase incremental training method, and a novel Top-N decoding strategy, facilitating controlled reward-guided generation using pretrained LLMs. To showcase the superior performance of ScaffoldGPT, we conduct evaluations on real-world viral and cancer-related datasets, comparing it against eight competing baselines. Our results demonstrate that ScaffoldGPT surpasses all baselines across the majority of performance metrics, underscoring its efficacy. Our work highlights ScaffoldGPT’s effectiveness in drug optimization, as evidenced by enhancements in various pharmaceutical properties. Currently, ScaffoldGPT is limited to handling molecules in SMILES format. We are looking into expanding ScaffoldGPT’s capabilities to accommodate a broader range of drug representation formats as our future direction.
Acknowledgements
This work is supported in part by the RadBio-AI project (DE-AC02-06CH11357), U.S. Department of Energy Office of Science, Office of Biological and Environment Research, the Improve project under contract (75N91019F00134, 75N91019D00024, 89233218CNA000001, DE-AC02-06-CH11357, DE-AC52-07NA27344, DE-AC05-00OR22725), the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration.
Impact Statement
This paper explores advancements in machine learning for developing language models aimed at drug optimization. While language models for drug optimization may have certain societal implications, we do not find any that must be specific emphasis here.
Appendix A Appendix
A.1 Pre-training and finetuning dataset
For pretrainig, We used the ZINC dataset, filtering for Standard, In-Stock, and Drug-Like molecules, resulting in approximately 11 million molecules. For preprocessing, we perform a few straightforward steps: 1. We first canonicalize smiles Chem.MolToSmiles(Chem.MolFromSmiles(mol),True). 2. We filter out molecules whose scaffold SMILES is an empty string. These preprocessing steps are also included in the huggingface data repo.
For finetuning, we utilize 1 million compounds from the ZINC15 dataset, docked to the 3CLPro protein (PDB ID: 7BQY) linked to SARS-CoV-2 and the RTCB protein (PDB ID: 4DWQ) associated with human cancer, as sourced from the latest Cancer and COVID dataset by Liu et al. [22], across all baselines.
A.2 Generation with finetuned model
The top five epochs with the highest historical average normalized reward (as detailed in Section 5.1) are selected. From these five epochs, the epoch with the highest product of validity and average normalized reward is chosen as the final model for generation.
With this epoch and corresponding weights, we apply the proposed decoding method (as described in section 4) for generation.
A.3 BPE Tokenization
The Byte Pair Encoding (BPE) algorithm involves the following steps:
  1. 1.
    Initialize the Vocabulary: Start with a base vocabulary consisting of all individual characters in the text corpus.
  2. 2.
    Count Frequencies: Count the frequency of all character pairs in the text.
  3. 3.
    Merge Most Frequent Pair: Identify the most frequent pair of characters and merge them into a single token. Add this new token to the vocabulary.
  4. 4.
    Update Text: Replace all occurrences of the most frequent pair with the new token in the text.
  5. 5.
    Repeat: Repeat the process of counting frequencies, merging pairs, and updating the text until the desired vocabulary size is reached or no more merges are possible.
By iteratively merging the most frequent pairs, BPE builds a robust vocabulary that captures common subword units, allowing for more efficient and flexible text representation.
A.4 Surrogate model
The surrogate model [44] is a simplified variant of a BERT-like transformer, extensively utilized in natural language processing. In this model, tokenized SMILES strings are inputted and then embedded with positional information. The resulting outputs are subsequently fed into a series of five transformer blocks, each comprising a multi-head attention layer (21 heads), a dropout layer, layer normalization with residual connection, and a feedforward network. This feedforward network consists of two dense layers followed by dropout and layer normalization with residual connection. Following the stack of transformer blocks, a final feedforward network is employed to generate the predicted docking score. The validation $r^{2}$ values are 0.842 for 3CLPro and 0.73 for the RTCB dataset.
A.5 Drug Optimization illustration on COVID benchmark
This is another example illustrating the effectiveness of ScaffoldGPT in enhancing the original molecule on the COVID benchmark. The results in Table 5 show that the drugs generated by ScaffoldGPT outperform the original drugs across all desired properties. Even though the original scaffold is altered and not present in the generated molecules, the similarity still demonstrates a decent level.
Original                        Scaffold improved 1 improved 2
Molecule [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE String c1ccc(C[N@H+]2CCC3(CC[NH+] (Cc4nccs4)CC3)C2)nc1 CC1(c2cccc(C(=O) Nc3ccccc3C)c2)CCCC1 Cc1ccc(CC2CCN (Cc3ccccn3)CC2)s1
Scaffold c1ccc(C[N@H+]2CCC3(CC[NH+] (Cc4nccs4)CC3)C2)nc1 - -
Docking ()($\downarrow$)( ↓ ) -9.748 -10.184 $\checkmark$ -10.187 $\checkmark$
Druglikeness ()($\uparrow$)( ↑ ) 0.839 0.840 $\checkmark$ 0.847 $\checkmark$
Synthesizability ()($\downarrow$)( ↓ ) 5.631 1.983 $\checkmark$ 2.199 $\checkmark$
Solubility ()($\uparrow$)( ↑ ) 0.192 5.079 $\checkmark$ 3.906 $\checkmark$
Similarity ()($\uparrow$)( ↑ ) - 0.335 0.563
Avg Norm Reward ()($\uparrow$)( ↑ ) 0.398 0.688 $\checkmark$ 0.694 $\checkmark$
Table 5: One molecule example from 3CLPro dataset, where scaffold and original are same. In this case the model tries to modify the scaffold, and the generated molecules does not contain scaffold. $\checkmark$ indicates improved property.
A.6 Binding sites of 3clpro and RTCB
(a) 3CLPro.
(b) RTCB.
Figure 2: The binding sites of proteins 3CLPro (PDB ID: 7BQY) (Left) and RTCB (PDB ID: 4DWQ) (Right). Binding sites are defined around the crystallized compound using Open Eye software.
A.7 Computing infrastructure and wall-time comparison
We trained our docking surrogate models using 4 nodes of the supercomputer where each node contains CPUs (64 cores) and 4 A100 GPU nodes [8]. The training time for each model was approximately 3 hours.
We conducted other experiments on a cluster that includes CPU nodes (approximately 280 cores) and GPU nodes (approximately 110 Nvidia GPUs, ranging from Titan X to A6000, set up mostly in 4- and 8-GPU configurations).
The pretraining process utilizes 8 GPUs, while APO and generation employs a single GPU. Both processes use either V100 or A100 GPUs. Based on the computing infrastructure, we obtained the wall-time comparison in Table 6 as follows.
Methods Total Run Time
Pretraining 24h
APO 27h
Top-N (One Generation) 17-20s
Table 6: Wall-time comparison between different methods.
A.8 Hyperparameters and architectures
Table 7 and 8 provides a list of hyperparameter settings we used for our experiments.
For APO finetuning and experimentation, 1280 molecules were selected from each of the RTCB and 3CLPro datasets, with docking scores ranging from -14 to -6. This range is based on [23].
Moreover, when computing the average normalized reward for the original molecule, in the absence of similarity considerations, we use weights of $0.25$ for docking, drug-likeness, synthesizability, and solubility, respectively.
Moreover, when the generated SMILES is invalid, indicating that the reward $R_{c}$ cannot be calculated, we have two options: either directly subtract the reward of the original SMILES (i.e., $-R_{c}(X)$), or consider the advantage preference as zero instead.
Parameter Value
Pretraining
Learning rate $5\times e^{-5}$
Batch size $24$
Optimizer Adam
# of Epochs for Training First Phase $10$
# of Epochs for Training Second Phase $10$
Model # of Params $124M$
Generation
N (Top-N) $1$
K (Number of possible next token) $16$
TopK $[10,15,20]$
TopP [$0.85$, $0.9$, $0.95$]
Table 7: Hyperparameters for pretraining and generation.
Parameter Value
Shared
# of Molecules Optimized $1280$
Learning rate $1\times 10^{-4}$
Optimizer Adam
# of Epochs for Training $100$
Batch size $64$
Best-of-N $[4,6,8]$
TopK $[10,15,20]$
TopP $[0.85,0.9,0.95]$
APO Objective Weight
Docking Score $0.2$
Druglikeliness $0.2$
Synthesizability $0.2$
Solubility $0.2$
Tamimoto Similarity $0.2$
APO Other
Fingerprint Size $1024$
Normalize Min/Max $[-10,10]$
Advantage preference with invalid generated SMILES
3CLPro $[0,-R_{c}(X)]$
RTCB $[0,-R_{c}(X)]$
Table 8: Hyperparameters for APO.